# Practical insights on the effect of different encodings, ansätze and measurements in quantum and hybrid convolutional neural networks

### v1

@@ III.3 Position of the quantum layer and effective entanglement
We found entanglement to be necessary for achieving top validation accuracy performance for the HQNN-Quanv architecture. In all cases where a non-entangling encoding and ansatz where used the performance was worse than the entangled alternatives. For the HQNN-Parallel, in some cases the non-entangled combination of ansatz and encoding was slightly better than some entangled versions. However, we would expect entangling combinations to perform better for higher qubit counts, as in our experiments each circuit only had access to $8/768$ features.
These observations should be taken with a grain of salt, given that for the hybrid models we only measured local observables and outcomes could be very different for other type of measurements.
III.4 Encoding, measurement and their interplay
For both hybrid models we found the encoding to be the most determining factor, FIG. 5, 6. Exceeding the effect of both ansatz and measurement, FIG. 11, 15.
We speculate this is due to several factors. First, the encoding is known to be a double edge sword: low complexity encodings don’t have enough expressive power Schuld et al. (2021) while excessively complex encodings lead to data concentration Li et al. (2023), which limits the distinguishability of encoded states. Secondly, we tried few ansätze compared to encodings. Perhaps, by adding a range of carefully selected options with varying complexities the impact would be more clear. Lastly, local measurements on Pauli $X/Y/Z$ may be too simple to meaningfully affect performance.
Amplitude encoding thrived only when the circuit processed a compressed global representation; it faltered whenever it was fed directly with small pixel windows, because normalising entire patches dilutes the very gradients that early layers should preserve.
@@ III.4 Encoding, measurement and their interplay
For hybrids, the measurements were significantly less decisive, which is logical as they are highly similar to each other and can be interconverted via single-qubit rotations. Therefore three Cartesian single-qubit bases produced practically identical scores, implying that the subsequent classical layers rotate whatever Bloch-vector projection they are given. In a purely quantum network, where no further classical post-processing is available until the final softmax, the simplest global measurement in the computational basis offered the clearest separation of classes, whereas multiqubit Pauli strings introduced unnecessary complexity without a matching gain. The study was not particularly exhaustive in this aspect, and given that these measurements are more distinct from one another than in the previous case, it is reasonable to expect increased variability among them.
@@
Interestingly, the fully quantum architectures used only amplitude encoding, yet a simple re-ordering of amplitudes before the first convolution layer delivered a noticeable lift. By arranging neighbouring pixels so that two-qubit gates act on semantically related amplitudes, the circuit effectively implements shaped kernels, mirroring how classical convolutions exploit spatial proximity.
III.5 Practical implications and open questions
@@
When validation accuracy is plotted against the number of trainable parameters, the classical baselines reveal the familiar pattern of saturating returns. The fully quantum models, however, achieve the same operating point while maintaining an order-of-magnitude smaller parameter budget. Although present-day simulation makes wall-clock training slower, the marked reduction in learnable weights hints at a tangible memory advantage for future quantum accelerators. The promising aspect here is the remarkably small size of the model and the significant results it can achieve with so few parameters. The accuracy-parameter ratio is substantially higher compared to classical counterparts.
@@
Several avenues remain open. First, the sweet-spot in entanglement may drift once realistic hardware noise is included, so repeating the sweep on early fault-tolerant devices would clarify the hardware-aware optimum. Moreover, analyzing the nonstabilizerness of the states generated by the different encodings and ansatze may shed light onto which combinations are best suited to provide value with respect to the classical counterpart. Efforts are already being made to study the behavior of such metrics in the context of PQCs Capecci et al. (2025). Second, data re-uploading layers—whose parameters modulate the encoding itself—could compensate for the shortcomings of an early quanvolution by learning a task-specific Fourier spectrum. Finally, a systematic survey of measurement strategies under realistic noise models is still lacking; understanding which observables remain informative when decoherence is present will be key to deploying quantum classifiers in practice.
IV Conclusion
@@
The study addresses the challenge of understanding how design choices in PQCs, such as data encoding, variational ansätze, and measurement bases, affect the performance of quantum and hybrid neural network architectures in image classification tasks using the EuroSAT dataset Helber et al. (2019).
@@
Hybrid models are most sensitive to data encoding strategies, with amplitude-based embeddings being incompatible with certain architectures (e.g., HQNN-Quanv) but effective in others (e.g., HQNN-Parallel). Variational ansätze and measurement bases have a secondary impact but the entangled ansätze, better in HQNN-Quanv than HQNN-Parallel. Purely quantum models, constrained to amplitude encoding, show performance dependence on the combination of measurement strategies and input ordering. These last models are considerably smaller than the current classics.
@@
As future work, it would be interesting to consider the aforementioned points; like adjusting hyperparameters or even exploring different image datasets. On the other hand, a thorough analysis of the parameter-accuracy relationship will be crucial to select a suitable model for implementation in quantum computers.
@@ Acknowledgments
The authors would like to express their gratitude for the collaborative efforts within the ARQA Cervera Technology Transfer Network Red ARQA (2024). This work is a result of the collective contributions and support from all members of this network, which has significantly enhanced our research outcomes.
@@
We would like to thank Andrés García Mangas and Marco Aguado Acevedo for his support in launching and executing the experiments using the CTIC Quantum Testbed (QUTE) platform García Mangas and Alonso (2025).
@@ A.1 The math behind
It is worth noting that with the same kind of reasoning we can see that a general unitary acting of three adjacent qubits can implement a kernel that mixes $82^{3}=$ 8 amplitudes of the state vector and, in general, we can create a kernel over $2^{K}$ pixels with a gate operating on $K$ adjacent qubits.
@@ A.2 Custom orderings
Figure 8: Diagram of how the pixels are ordered in the amplitudes of the state vector of the $N$-qubit circuit. For each ordering, the picture on the left (right) represents the which pixels the first (second) gate is mixing. Pixels shaded with the same color are mixed together.
@@ Appendix B HQNN Parallel Hybrid model
Figure 10: Training and validation accuracy and loss for the best hybrid model compared with the classic one.
@@ Appendix C HQNN-Quanv model results
Figure 14: Train and validation accuracy and loss for the best hybrid model compared with the classic one.
@@ Appendix D Pure quantum models results
Figure 18: Training and validation accuracy and loss for the best purely quantum model for 4 and 10 classes.
@@ Appendix E QCNN and SEQNN, full results overview
Paulis 63,75 42,1 300
Flatten histogram 59,75 - 300
PauliZ 68,5 - 293
Paulis 63,5 42,5 300
Two kernel Random histogram 60,5 - 184
PauliZ 60,25 -
Paulis 34,25 31,3
Squared histogram 60,75 - 184
PauliZ 64,75 -
Paulis 35,75 31,9
FC (Simplified Two Design) Random histogram 60,25 - 190
PauliZ 66 36,8
Paulis 53,25 28,2
Squared histogram 62,75 - 190
PauliZ 62,25 38,5
Paulis 55,5 29,1
@@
QCNN Random histogram 62,25 - 326
PauliZ 66,75 - 373
Paulis 63,5 41,5 316
Flatten histogram 59,25 - 326
PauliZ 63,5 - 373
Paulis 64,5 41 316
Two kernel Random histogram 53,75 - 310
PauliZ 57,75 -
@@
Paulis 38,75 37,1
FC (Simplified Two Design) Random histogram 60,25 - 352
PauliZ 60,75 39,3
Paulis 54 29,8
Squared histogram 62,5 - 352
PauliZ 60,5 40,3
Paulis 53,25 30,5

### v2

@@ V.3 Position of the quantum layer and effective entanglement
We found entanglement to be necessary for achieving top validation accuracy performance for the HQNN-Quanv architecture. In all cases where a non-entangling encoding and ansatz were used, the performance was worse than the entangled alternatives. For the HQNN-Parallel, in some cases the non-entangled combination of ansatz and encoding was slightly better than some entangled versions. However, we would expect entangling combinations to perform better for higher qubit counts, as in our experiments each circuit only had access to $8/768$ features.
These observations should be taken with a grain of salt, given that for the hybrid models we only measured local observables, and outcomes could be very different for other types of measurements.
V.4 Encoding, measurement, and their interplay
For both hybrid models, we found the encoding to be the most determining factor, FIG. 8, 9. Exceeding the effect of both ansatz and measurement, FIG. 14, 18.
We speculate this is due to several factors. First, the encoding is known to be a double-edged sword: low-complexity encodings don’t have enough expressive power schuldEffectDataEncoding2021 while excessively complex encodings lead to data concentration liConcentrationDataEncoding, which limits the distinguishability of encoded states. Secondly, we tried few ansätze compared to encodings. Perhaps, by adding a range of carefully selected options with varying complexities, the impact would be more clear. Lastly, local measurements on Pauli $X/Y/Z$ may be too simple to meaningfully affect performance.
Amplitude encoding thrived only when the circuit processed a compressed global representation; it faltered whenever it was fed directly with small pixel windows, because normalizing entire patches dilutes the very gradients that early layers should preserve.
@@ V.4 Encoding, measurement, and their interplay
For hybrids, the measurements were significantly less decisive, which is logical as they are highly similar to each other and can be interconverted via single-qubit rotations. Therefore, three Cartesian single-qubit bases produced practically identical scores, implying that the subsequent classical layers rotate whatever Bloch-vector projection they are given. In a purely quantum network, where no further classical post-processing is available until the final softmax, the simplest global measurement in the computational basis offered the clearest separation of classes, whereas multiqubit Pauli strings introduced unnecessary complexity without a matching gain. The study was not particularly exhaustive in this aspect, and given that these measurements are more distinct from one another than in the previous case, it is reasonable to expect increased variability among them.
@@
Interestingly, the fully quantum architectures used only amplitude encoding, yet a simple reordering of amplitudes before the first convolution layer delivered a noticeable lift. By arranging neighboring pixels so that two-qubit gates act on semantically related amplitudes, the circuit effectively implements shaped kernels, mirroring how classical convolutions exploit spatial proximity.
V.5 Practical implications and open questions
@@
When validation accuracy is plotted against the number of trainable parameters, the classical baselines reveal the familiar pattern of saturating returns. The fully quantum models, however, achieve the same operating point while maintaining an order-of-magnitude smaller parameter budget. Although present-day simulation makes wall-clock training slower, the marked reduction in learnable weights hints at a tangible memory advantage for future quantum accelerators. The promising aspect here is the remarkably small size of the model and the significant results it can achieve with so few parameters. The accuracy-parameter ratio is substantially higher in the QCNN compared to all others as we can see in TABLE 2.
Model No. of params Accuracy (%) Accuracy- Param $\cdot$ 100
Classical ”Parallel” $903598$ $67.1$ $0.007$
Classical ”Quanv” $273706$ $71.5$ $0.026$
HQNN-Parallel $15658$ $74.8$ $0.477$
HQNN-Quanv 930826 $65.9$ $0.007$
QCNN $300$ $42.5$ $15$
Table 2: Comparison of parameter count and accuracy-parameter ratio across models. For all hybrid-quantum models, the optimal configuration was selected. HQNN-Parallel: Amplitude-No Entanglement; HQNN-Quanv: AngleX-Full Entanglement; QCNN-Flatten-Paulis.
@@
Several avenues remain open. First, the sweet-spot in entanglement may drift once realistic hardware noise is included, so repeating the sweep on early fault-tolerant devices would clarify the hardware-aware optimum. Moreover, analyzing the non-stabilizerness of the states generated by the different encodings and ansatze may shed light on which combinations are best suited to provide value regarding the classical counterpart. Efforts are already being made to study the behavior of such metrics in the context of PQCs capecci2025role. Second, data re-uploading layers—whose parameters modulate the encoding itself—could compensate for the shortcomings of an early quanvolution by learning a task-specific Fourier spectrum. Finally, a systematic survey of measurement strategies under realistic noise models is still lacking; understanding which observables remain informative when decoherence is present will be key to deploying quantum classifiers in practice.
It is crucial to interpret these results within the context of ideal, noiseless simulations. While our analysis identifies a performance hierarchy where entangled ansätze generally outperform non-entangled variants (particularly in the HQNN-Quanv architecture), this advantage may be compromised in a real-world NISQ environment. On physical hardware, the introduction of entanglement increases circuit depth and gate count, thereby accumulating decoherence and operational errors. Consequently, a trade-off emerges where the theoretical expressibility gained by entanglement must be weighed against the hardware-induced noise. Future work must rigorously validate these findings under realistic noise models to determine the threshold at which the benefits of entanglement are negated by gate errors.
VI Conclusion
@@
The study addresses the challenge of understanding how design choices in PQCs, such as data encoding, variational ansätze, and measurement bases, affect the performance of quantum and hybrid neural network architectures in image classification tasks using the EuroSAT dataset Helber2019.
@@
Hybrid models are most sensitive to data encoding strategies, with amplitude-based embeddings being incompatible with certain architectures (e.g., HQNN-Quanv) but effective in others (e.g., HQNN-Parallel). Variational ansätze and measurement bases have a secondary impact, but the entangled ansätze performed better in HQNN-Quanv than HQNN-Parallel. Purely quantum models, constrained to amplitude encoding, show performance dependence on the combination of measurement strategies and input ordering. The latter models are considerably smaller than the current classical models.
@@
As future work, it would be interesting to consider the aforementioned points, like adjusting hyperparameters or even exploring different image datasets. On the other hand, a thorough analysis of the parameter-accuracy relationship will be crucial to select a suitable model for implementation in quantum computers.
@@ Acknowledgments
The authors would like to express their gratitude for the collaborative efforts within the ARQA Cervera Technology Transfer Network ARQA. This work is a result of the collective contributions and support from all members of this network, which has significantly enhanced our research outcomes.
@@
We would like to thank Andrés García Mangas and Marco Aguado Acevedo for his support in launching and executing the experiments using the CTIC Quantum Testbed (QUTE) platform qute.
@@ A.1 The math behind
It is worth noting that with the same kind of reasoning we can see that a general unitary acting of three adjacent qubits can implement a kernel that mixes $2^{3}=8$ amplitudes of the state vector and, in general, we can create a kernel over $2^{K}$ pixels with a gate operating on $K$ adjacent qubits.
@@ A.2 Custom orderings
Figure 11: Diagram of how the pixels are ordered in the amplitudes of the state vector of the NN-qubit circuit. For each ordering, the picture on the left (right) represents which pixels the first (second) gate is mixing. Pixels shaded with the same color are mixed together.
@@ Appendix B HQNN Parallel Hybrid model
Figure 13: Training and validation accuracy and loss for the best HQNN-Parallel model (Amplitude-No entanglement) compared with the classical one.
@@ Appendix C HQNN-Quanv model results
Figure 17: Train and validation accuracy and loss for the best HQNN-Quanv model (AngleX-Full entanglement) compared with the classical one.
@@ Appendix D Pure quantum models results
Figure 21: Training and validation accuracy and loss for the best purely quantum model for 4 (QCNN-Flatten-PauliZ) and 10 (QCNN-Flatten-Paulis) classes.
@@ Appendix E QCNN and SEQNN, full results overview
Paulis 63.75 42.1 300
Flatten histogram 59.75 - 300
PauliZ 68.5 - 293
Paulis 63.5 42.5 300
Two kernel Random histogram 60.5 - 184
PauliZ 60.25 -
Paulis 34.25 31.3
Squared histogram 60.75 - 184
PauliZ 64.75 -
Paulis 35.75 31.9
FC (Simplified Two Design) Random histogram 60.25 - 190
PauliZ 66 36.8
Paulis 53.25 28.2
Squared histogram 62.75 - 190
PauliZ 62.25 38.5
Paulis 55.5 29.1
@@
QCNN Random histogram 62.25 - 326
PauliZ 66.75 - 373
Paulis 63.5 41.5 316
Flatten histogram 59.25 - 326
PauliZ 63.5 - 373
Paulis 64.5 41 316
Two kernel Random histogram 53.75 - 310
PauliZ 57.75 -
@@
Paulis 38.75 37.1
FC (Simplified Two Design) Random histogram 60.25 - 352
PauliZ 60.75 39.3
Paulis 54 29.8
Squared histogram 62.5 - 352
PauliZ 60.5 40.3
Paulis 53.25 30.5
@@
Appendix F Expressibility of QAOA encoding
In this section we calculate an estimate of the state frame potential for the $QAOA$ encoding for the three possible local fields. The state frame potential has been used before as a measure of expressibility by comparing with that of the ensemble of Haar distributed states sim2019expressibility. We compare the first two moments, $\mathcal{F}^{(1)}$ and $\mathcal{F}^{(2)}$ of the frame potential normalized by the value for Haar states for the three local fields. For that, we calculate the measure over several fixed random inputs. The lower the measure, the closer the distribution of states generated by the encoding circuits is to the Haar distribution; hence, more expressive. Results are conveyed in TABLE 5 and TABLE 6.
Local field $\mathbf{\mathcal{F}^{(1)}/\mathcal{F}^{(1)}_{Haar}}$ $\mathbf{\mathcal{F}^{(2)}/\mathcal{F}^{(2)}_{Haar}}$
X $10.02(0.23)$ $55.6(2.5)$
Y $9.67(0.23)$ $52.6(2.3)$
Z $12.5(1.0)$ $96(14)$
Table 5: Results for the state frame potential for $4$ qubits.
Local field $\mathbf{\mathcal{F}^{(1)}/\mathcal{F}^{(1)}_{Haar}}$ $\mathbf{\mathcal{F}^{(2)}/\mathcal{F}^{(2)}_{Haar}}$
X $100.0(4.6)$ $5550(460)$
Y $96.1(3.4)$ $5220(360)$
Z $154(21)$ $12800(3400)$
Table 6: Results for the state frame potential for $8$ qubits.